OpenStudy (anonymous):

Solve for x=4 y=2x+3

OpenStudy (anonymous):

y = 2*4 + 3 y = 8 + 3 y = 11

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

y = 2(4) + 3 y = 8 + 3 y = 11

OpenStudy (anonymous):

Thanks!